OpenStudy (anonymous):

Use cross products to demonstrate whether or not the two ratios are equivalent: 4/6 and 14/21 .

OpenStudy (amistre64):

\[\frac{4}{6}=\frac{14}{21}\] \[\frac{\frac{4}{6}}{\frac{14}{21}}=\frac{\frac{14}{21}}{\frac{14}{21}}\] \[\frac{\frac{4}{6}}{\frac{14}{21}}=1\] \[\frac{\frac{21}{14}\frac{4}{6}}{\frac{21}{14}\frac{14}{21}}=1\] \[\frac{\frac{21}{14}\frac{4}{6}}{1}=1\] \[Does~\frac{21}{14}\frac{4}{6}=1~~?\]

OpenStudy (amistre64):

or tp convert it into cross priduct form \[\frac{21}{14}\frac{4}{6}=1\] \[(14~6)\frac{21}{14}\frac{4}{6}=1(14~6)\] \[21\times4=14\times6\]